OpenStudy (anonymous):

Solve tan(x)(tan(x)-1)=0 a. x=pi/3±2pi(n),x=3pi/4±2pin b. x=pi/2±2pi(n),x=3pi/2±2pi(n),x=±2pi(n) c. x=±pi(n),x=pi/2±2pi(n) d. x=±n(pi),x=pi/4±n(pi)

OpenStudy (anonymous):

tan(x)(tan(x)-1)=0 this only happens when tan x = 1 ( sin(x)=cos(x) )

OpenStudy (anonymous):

and sin x = cos x only in the first quarter the shaded area 

OpenStudy (anonymous):